OpenStudy (fanduekisses):

CALCULUS DERIVATIVE help: At a sand and gravel plant, sand is falling off a conveyor and onto a conical pile at a rate of 10 cubi feet per minute. The diameter of the base of the one is approx three times the altitude At what rate is the height of the pile changing when the pile is 15 feet high?

OpenStudy (fanduekisses):

Ok so basically, I'm looking for the derivative of the height right?

OpenStudy (fanduekisses):

So I use the formula to find the volume of a cone?

OpenStudy (wmj259):

Rate of Change. Correct?

OpenStudy (fanduekisses):

yes, of the height.

OpenStudy (surry99):

You will need an expression for the volume of a cone and will have to use the chain rule to obtain a relationship between dv/dt and dh/dt

OpenStudy (fanduekisses):

ok so v=pi*r^2*h/3

OpenStudy (fanduekisses):

dv/dt=pi*2r*dr/dt* -3h^-2 * dh/dt

OpenStudy (surry99):

Your equation for Volume is correct but the derivative is not. Also note the problem told you something about the diameter of the base of the cone. You will need to use that...do you see why?

OpenStudy (fanduekisses):

oooooh ok

OpenStudy (fanduekisses):

so the diameter is 3x the height?

OpenStudy (surry99):

V = 1/3*pi*r^2*h agreed

OpenStudy (surry99):

Now the diameter is 1/3 of the height...but the Volume formula does not have the diameter in it...how do you introduce it?

OpenStudy (fanduekisses):

2*r

OpenStudy (surry99):

ok so r = d/2 and d=h/3 substitute these into your formaula for the Volume and lets see what you get

OpenStudy (anonymous):

2r=3h r=?

OpenStudy (fanduekisses):

r=3h/2

OpenStudy (surry99):

So what does the Volume formula simplify to?

OpenStudy (fanduekisses):

but wait, I'm confused now...hmmm after substituting I get v= 1/3*pi*3h/2

OpenStudy (surry99):

Do it one step at a time.

OpenStudy (anonymous):

\[V=\frac{ 1 }{ 3 } \pi r^2 h=\frac{ \pi }{ 3 }\left( \frac{ 3h }{ 2 } \right)^2h=\frac{ 3 \pi }{ 4 }h^3\] dV/dt=?

OpenStudy (fanduekisses):

ok so the volume of a cone is π*r^2*h/3 The diameter is 3x the height. and radius is d/2

OpenStudy (surry99):

So V = what?

OpenStudy (fanduekisses):

but how do you get 3h/2? because After plugging in h/3 in r=d/2, I get 2h/3 instead..

OpenStudy (surry99):

Look at what surithayer got...do you get the same?

OpenStudy (surry99):

My mistake d = 3h

OpenStudy (fanduekisses):

yeah because it's 3 times

OpenStudy (fanduekisses):

ok

OpenStudy (fanduekisses):

and where did the 1/3 come from?

OpenStudy (surry99):

that was my mistake ...sorry.

OpenStudy (fanduekisses):

the 1/3 ?

OpenStudy (fanduekisses):

When I google the formula for the volume of a cone I get v=pi*r^2*h/3

OpenStudy (surry99):

ok V = 1/3 pi* r^2 * h that is the volume of a cone...so the 1/3 comes from this formula

OpenStudy (surry99):

yes..see the 1/3 at the end?

OpenStudy (fanduekisses):

I thought the 3 is only under the h

OpenStudy (surry99):

So V = 1/3*pi*r^2*h which is the same as pi*r^2*h/3

OpenStudy (fanduekisses):

oh ok :)

OpenStudy (fanduekisses):

So now that v=3pi/4 *h^3, can I take the derivative?

OpenStudy (surry99):

Yes...but take it carefully using the chain rule. Lets see what you get.

OpenStudy (fanduekisses):

dv/dt = 3pi/4 *3h^2 *dh/dt

OpenStudy (surry99):

Yes...now substitute and solve for dh/dt

OpenStudy (fanduekisses):

ok ^_^

OpenStudy (surry99):

Get an answer?

OpenStudy (fanduekisses):

Ok so dh/dt = 8/405 pi :D

OpenStudy (fanduekisses):

yay thanks so much ^_^

OpenStudy (surry99):

I got .00629 ft/min wheras your answer works out to .062...just double check

OpenStudy (surry99):

your welcome ...